# The Efficiency Spectrum of Large Language Models: An Algorithmic Survey

### v1

@@ 6.2.3. Parallelism-Based Techniques
$\bullet$ Pipeline Parallelism (Jia et al., 2019; Huang et al., 2019; Narayanan et al., 2019, 2021a, 2021b; Yang et al., 2021; Kosson et al., 2021; Kim et al., 2023b; Li et al., [n. d.]), on the other hand, is a form of inter-layer model parallelism. It involves splitting the layers of a model across multiple accelerators in a pipeline configuration. Each accelerator is responsible for computing a different layer and then passing its output to the next, akin to an assembly line. This setup allows for sequential yet concurrent processing of both forward and backward passes. While one segment of the model processes one part of the data, other segments can work on different parts simultaneously. The key to the efficiency of PP lies in its ability to keep all parts of the model active and productive, though it requires careful scheduling to minimize idle time on the GPUs.
Gpipe (Huang et al., 2019), one of the earliest proposals for PP, combines PP with mini-batch splitting. It segments a large model across multiple GPUs and processes input mini-batches as smaller micro-batches. This approach allows for the efficient training of significantly large models. GPipe also uses a strategy called rematerialization (Chen et al., 2016) to reduce memory usage by recalculating activations during backward propagation instead of storing them. However, GPipe still faces memory inefficiencies due to the need to cache activations during backward computations. PipeDream (Harlap et al., 2018; Narayanan et al., 2019) further refines the PP approach with its One Forward pass followed by One Backward pass (1F1B) strategy. This method allows for the immediate backward propagation of a micro-batch following its forward pass, enabling earlier stages of the pipeline to start their backward computations sooner. PipeDream also employs asynchronous gradient updates using different versions of weights and optimizes memory allocation across the pipeline stages for more efficient processing. BPipe (Kim et al., 2023b) introduces a technique to balance memory usage throughout the pipeline. It leverages idle memory in later stages to support earlier stages, significantly speeding up the training process for large models like GPT-3 $96$B and GPT-3 $134$B by $1.25$ to $2.17$ times compared to Megatron-LM. TeraPipe (Li et al., [n. d.]) addresses the challenge of training large models with extensive sequence lengths, which can lead to smaller mini-batches and increased idle time in the pipeline. In transformer architectures, some layers’ calculations don’t depend on future hidden states. TeraPipe uses this fact to enable parallel processing by dividing the input sequence. It uses dynamic programming to effectively split the sequence at the best points across tokens, improving the efficiency of parallel processing.
@@
To streamline this process, automated parallelization solutions are being developed. These solutions aim to speed up model deployment and ensure adaptability across different models. As models and computing clusters grow, the complexity of parallelism configurations also increases. Tofu (Wang et al., [n. d.]) tackles this challenge with a dynamic programming algorithm that optimizes dataflow graph partitioning. Dapple (Fan et al., 2021) focuses on minimizing pipeline latency through optimal partitioning strategies. However, these solutions are currently limited to combining data parallelism with only one model parallelism type, due to the complex interactions between different parallelism strategies. Alpa (Zheng et al., 2022) takes a more comprehensive approach by organizing data, model, and pipeline parallelism into a hierarchical structure. It uses integer linear programming for model parallelism plans and dynamic programming for pipeline parallelism plans. Alpa is comparable to specialized systems like Megatron-LM in training GPT models, demonstrating its effectiveness in handling complex parallelization challenges.
@@ 6.3.1. Parameter-Efficient Fine-Tuning (PEFT)
Traditional methods required manual design of prompt templates and verbalizers, which often resulted in sensitive and varying efficacy. However, recent advancements in prompt learning have led to the automation and optimization of prompt construction. AutoPrompt (Shin et al., 2020) introduces a gradient-based approach to automate the search for effective templates. LM-BFF (Gao et al., 2020) offers a more efficient solution for automated prompt generation by searching for label words and using a T5-based template generation method in the discrete prompt space. To tackle the challenges of discrete optimization, Prefix-Tuning (Li and Liang, 2021) recommends parameterized prompts, where only the prompt is fine-tuned while the LLM remains unaltered. P-tuning (Liu et al., 2023d) breaks away from the conventional constraint that templates must be composed of natural language, transforming template construction into a continuous parameter optimization challenge. CP-tuning (Xu et al., 2023) advocates the use of contrastive learning to automatically learn the distribution of embeddings, serving as a substitute for manual verbalizer design. UPT (Wang et al., 2022c) introduces the Prompt-Options-Verbalizer paradigm, facilitating joint prompt learning across diverse NLP tasks and encouraging LLMs to acquire task-invariant prompting knowledge.
@@ 7.2. Pruning
Unstructured Pruning. Unstructured pruning yields fine-grained sparsity wherein zero elements are randomly distributed across the trainable parameters (Jaiswal et al., 2023; SAUCE, [n. d.]; Shao et al., 2023; Zhang et al., 2023e; Kurtic et al., 2023; Xia et al., 2023b; Sun et al., 2023b; Chen et al., 2020a; Ding et al., 2021, 2022; Chen et al., 2020b). These unstructured pruning methods show that LLMs can be pruned to at least 50% sparsity in one-shot, with(out) retraining, at minimal loss of accuracy. Meanwhile, larger LLMs are more compressible in the manner of unstructured pruning. Although unstructured pruning could bring theoretical inference speedup, the speedup is not easily reflected in reality due to the contiguity issue during sparse operations. Accelerating a DNN with high fine-grained sparsity typically requires the supports of specialized designed software and hardware. There exist computing libraries such as FSCNN (Ji and Chen, 2022) which could outperform standard DNN runtime under sufficiently high unstructured sparsity, yet largely not extended to transformer architectures.
@@ 7.3. Knowledge Distillation
The concept of knowledge distillation involves utilizing supervisory signals from a large, more capable ‘teacher’ model to train a compact ‘student’ model. This approach often results in the student model surpassing the performance of a similarly sized model that was trained without such guidance (Hinton et al., 2015). Knowledge distillation can be largely categorized into response-based, feature-based, and relation-based knowledge distillation (Gou et al., 2021). Response-based knowledge focuses on the final output layer of the teacher model. The hypothesis is that the student model will learn to mimic the predictions of the teacher model (Hinton et al., 2015; Ba and Caruana, 2014). A trained teacher model also captures feature-based knowledge of the data in its intermediate layers, which is especially pertinent for deep neural networks (Romero et al., 2014; Zagoruyko and Komodakis, 2016). Knowledge that captures the relationship between feature maps can also be used to train a student model, referring as relation-based (Yim et al., 2017). Initial research in NLP domain primarily concentrated on the distillation of task-specific models (Kim and Rush, 2016). Later on, more studies have shifted their focus towards distilling pre-trained models, which can subsequently be fine-tuned for specialized downstream tasks (Sanh et al., 2019; Liu et al., 2020b; Jiao et al., 2019). Recently, there emerge distillation methods for LLMs (Zhang et al., 2023d; Li et al., 2023b; Hsieh et al., 2023; Wang et al., 2023a; Wu et al., 2023b; Chen et al., 2023b; Anand et al., 2023; Gu et al., 2023; Jiang et al., 2023; Chen et al., 2023f; Sahu et al., 2023). One main focus of the current knowledge distillation methods on LLMs lies in how to generate and utilize challenging (instructed) samples (Wu et al., 2023b; Chen et al., 2023b; Jiang et al., 2023) to more effectively transfer the knowledge from teacher to student model, which has some overlapping to the data undersampling methods in Section 4. Chain-of-thought prompting is commonly used in distillation approaches (Li et al., 2023b; Hsieh et al., 2023) to accomplish the data generations.

### v2

@@ 6.2.3. Parallelism-Based Techniques
$\bullet$ Pipeline Parallelism (Jia et al., 2019; Huang et al., 2019; Narayanan et al., 2019, 2021a, 2021b; Yang et al., 2021; Kosson et al., 2021; Kim et al., 2023b), on the other hand, is a form of inter-layer model parallelism. It involves splitting the layers of a model across multiple accelerators in a pipeline configuration. Each accelerator is responsible for computing a different layer and then passing its output to the next, akin to an assembly line. This setup allows for sequential yet concurrent processing of both forward and backward passes. While one segment of the model processes one part of the data, other segments can work on different parts simultaneously. The key to the efficiency of PP lies in its ability to keep all parts of the model active and productive, though it requires careful scheduling to minimize idle time on the GPUs.
Gpipe (Huang et al., 2019), one of the earliest proposals for PP, combines PP with mini-batch splitting. It segments a large model across multiple GPUs and processes input mini-batches as smaller micro-batches. This approach allows for the efficient training of significantly large models. GPipe also uses a strategy called rematerialization (Chen et al., 2016) to reduce memory usage by recalculating activations during backward propagation instead of storing them. However, GPipe still faces memory inefficiencies due to the need to cache activations during backward computations. PipeDream (Harlap et al., 2018; Narayanan et al., 2019) further refines the PP approach with its One Forward pass followed by One Backward pass (1F1B) strategy. This method allows for the immediate backward propagation of a micro-batch following its forward pass, enabling earlier stages of the pipeline to start their backward computations sooner. PipeDream also employs asynchronous gradient updates using different versions of weights and optimizes memory allocation across the pipeline stages for more efficient processing. BPipe (Kim et al., 2023b) introduces a technique to balance memory usage throughout the pipeline. It leverages idle memory in later stages to support earlier stages, significantly speeding up the training process for large models like GPT-3 $96$B and GPT-3 $134$B by $1.25$ to $2.17$ times compared to Megatron-LM. TeraPipe (literapipe) addresses the challenge of training large models with extensive sequence lengths, which can lead to smaller mini-batches and increased idle time in the pipeline. In transformer architectures, some layers’ calculations don’t depend on future hidden states. TeraPipe uses this fact to enable parallel processing by dividing the input sequence. It uses dynamic programming to effectively split the sequence at the best points across tokens, improving the efficiency of parallel processing.
@@
To streamline this process, automated parallelization solutions are being developed. These solutions aim to speed up model deployment and ensure adaptability across different models. As models and computing clusters grow, the complexity of parallelism configurations also increases. Tofu (Wang et al., 2018) tackles this challenge with a dynamic programming algorithm that optimizes dataflow graph partitioning. Dapple (Fan et al., 2021) focuses on minimizing pipeline latency through optimal partitioning strategies. However, these solutions are currently limited to combining data parallelism with only one model parallelism type, due to the complex interactions between different parallelism strategies. Alpa (Zheng et al., 2022) takes a more comprehensive approach by organizing data, model, and pipeline parallelism into a hierarchical structure. It uses integer linear programming for model parallelism plans and dynamic programming for pipeline parallelism plans. Alpa is comparable to specialized systems like Megatron-LM in training GPT models, demonstrating its effectiveness in handling complex parallelization challenges.
@@ 6.3.1. Parameter-Efficient Fine-Tuning (PEFT)
Traditional methods required manual design of prompt templates and verbalizers, which often resulted in sensitive and varying efficacy. However, recent advancements in prompt learning have led to the automation and optimization of prompt construction. AutoPrompt (Shin et al., 2020) introduces a gradient-based approach to automate the search for effective templates. LM-BFF (Gao et al., 2020) offers a more efficient solution for automated prompt generation by searching for label words and using a T5-based template generation method in the discrete prompt space. To tackle the challenges of discrete optimization, Prefix-Tuning (Li and Liang, 2021) recommends parameterized prompts, where only the prompt is fine-tuned while the LLM remains unaltered. P-tuning (Liu et al., 2023e) breaks away from the conventional constraint that templates must be composed of natural language, transforming template construction into a continuous parameter optimization challenge. CP-tuning (Xu et al., 2023) advocates the use of contrastive learning to automatically learn the distribution of embeddings, serving as a substitute for manual verbalizer design. UPT (Wang et al., 2022c) introduces the Prompt-Options-Verbalizer paradigm, facilitating joint prompt learning across diverse NLP tasks and encouraging LLMs to acquire task-invariant prompting knowledge.
@@ 7.2. Pruning
Unstructured Pruning. Unstructured pruning yields fine-grained sparsity wherein zero elements are randomly distributed across the trainable parameters (Jaiswal et al., 2023; SAUCE, 2023; Shao et al., 2023; Zhang et al., 2023e; Kurtic et al., 2023; Xia et al., 2023b; Sun et al., 2023b; Chen et al., 2020b, a; Ding et al., 2021, 2022). These unstructured pruning methods show that LLMs can be pruned to at least 50% sparsity in one-shot, with(out) retraining, at minimal loss of accuracy. Meanwhile, larger LLMs are more compressible in the manner of unstructured pruning. Although unstructured pruning could bring theoretical inference speedup, the speedup is not easily reflected in reality due to the contiguity issue during sparse operations. Accelerating a DNN with high fine-grained sparsity typically requires the supports of specialized designed software and hardware. There exist computing libraries such as FSCNN (Ji and Chen, 2022) which could outperform standard DNN runtime under sufficiently high unstructured sparsity, yet largely not extended to transformer architectures.
@@ 7.3. Knowledge Distillation
The concept of knowledge distillation involves utilizing supervisory signals from a large, more capable ‘teacher’ model to train a compact ‘student’ model. This approach often results in the student model surpassing the performance of a similarly sized model that was trained without such guidance (Hinton et al., 2015). Knowledge distillation can be largely categorized into response-based, feature-based, and relation-based knowledge distillation (Gou et al., 2021). Response-based knowledge focuses on the final output layer of the teacher model. The hypothesis is that the student model will learn to mimic the predictions of the teacher model (Hinton et al., 2015; Ba and Caruana, 2014). A trained teacher model also captures feature-based knowledge of the data in its intermediate layers, which is especially pertinent for deep neural networks (Romero et al., 2014; Zagoruyko and Komodakis, 2016). Knowledge that captures the relationship between feature maps can also be used to train a student model, referring as relation-based (Yim et al., 2017). Initial research in NLP domain primarily concentrated on the distillation of task-specific models (Kim and Rush, 2016). Later on, more studies have shifted their focus towards distilling pre-trained models, which can subsequently be fine-tuned for specialized downstream tasks (Sanh et al., 2019; Liu et al., 2020b; Jiao et al., 2019). Recently, there emerge distillation methods for LLMs (Zhang et al., 2023d; Li et al., 2023b; Hsieh et al., 2023; Wang et al., 2023a; Wu et al., 2023b; Chen et al., 2023b; Anand et al., 2023; Gu et al., 2023; Jiang et al., 2023; Chen et al., 2023f; Sahu et al., 2023). One main focus of the current knowledge distillation methods on LLMs lies in how to generate and utilize challenging (instructed) samples (Wu et al., 2023b; Chen et al., 2023b; Jiang et al., 2023) to more effectively transfer the knowledge from teacher to student model, which has some overlapping to the data undersampling methods in Section 4. Chain-of-thought prompting is commonly used in distillation approaches (Li et al., 2023b; Hsieh et al., 2023) to accomplish the data generations. However, current knowledge distillation methods for LLMs suffers from the absence of a standardized objective function, which is typically task specific and requires tuning efforts. Furthermore, the recent strategy of employing student-generated outputs to mitigate the discrepancies between training and inference has led to a substantial increase in computational costs. Addressing these challenges, DistiLLM (Ko et al., 2024) has been recently introduced, consisting of two key components. (i) A skew Kullback-Leibler (KL) divergence loss with moderate skew value, which offers stable gradients and lower approximation errors compared to other alternatives, including the generalized Jensen–Shannon divergence. (ii) An adaptive off-policy technique, designed to improve the efficiency and effectiveness of using student-generated outputs.